# Active Learning of Deep Neural Networks via Gradient-Free Cutting Planes

### v1

@@ 1 Introduction
2 Prior Work
Cutting-Plane Method.
Active Learning.
With the fast growth of current deep learning model sizes, more and more data is in need for training an effective deep NN model. Compared to just feeding the model with a set of randomly-selected training samples, how to select the most informative data for each training iteration becomes a valuable question. Efficient data sampling algorithm is thus increasingly important, especially for current RL-based language model training scheme. With the increasing power of current pretrained models, researchers found that since most of daily questions can be addressed correctly by the model already, these questions are valueless for further boosting the language models’ capacity. Therefore, actively selecting questions that LLM cannot address correctly is key to current LLM training (Lightman et al., 2023). Based on different learning settings, active learning strategies can be divided into: stream-based selective sampling used when the data is generated continuously (Woodward & Finn, 2017); pool-based sampling used when a pool of unlabeled data is presented (Gal et al., 2017b); query synthesis methods used when new samples can be generated for labeling. Based on different information measuring schemes, active learning algorithms can be further divided into uncertainty sampling (Lewis & Gale, 1994b) which selects data samples to reduce prediction uncertainty, query-by-committee sampling which involves multiple models for data selection, diversity-weighted method which selects the most diverse data sample, and expected error reduction method (Mussmann et al., 2022) which selects samples to best reduce models’ expected prediction error.
Convex NN.
@@ 5 Key Observation: Training ReLU NNs for Binary Classification is Linear Programming
To break this bottleneck and extend the cutting-plane-based learning method to nonlinear models, we observe that training a ReLU network for binary classification can be formulated as a linear programming problem. This insight is crucial for extending the learning scheme in (Louche & Ralaivola, 2015) to more complex models. We now develop our core idea of reframing binary classification with ReLU models as linear programs. For clarity, we present our results in two theorems: one for two-layer ReLU networks and another for ReLU networks of arbitrary depth. We focus on the two-layer case in the main paper for detailed discussion, deferring the more abstract general case to Appendix D.2. Since the general case is an extension of the two-layer model, focusing on the two-layer case should provide a clearer understanding of the core concepts.
@@
Therefore, if we iterate over all $D_{i}\in\mathcal{D}$, we are guaranteed to reach each ${K}_{i}$. The equivalence in Theorem 5.1 holds in the sense that, which we prove rigorously in Appendix D.1, whenever there is solution $W_{1},W_{2}$ to Problem (2), there is always solution $\{u_{i},u_{i}^{\prime}\}$ to problem (3) and vice versa. Moreover, when an optimal solution $\{u_{i}^{\star},u_{i}^{\prime\star}\}$ to problem (3) has been found, we can explicitly create an optimal solution $\{W_{1}^{\star},W_{2}^{\star}\}$ to Problem (2) from value of $\{u_{i}^{\star},u_{i}^{\prime\star}\}$, see Appendix D.1 for details. Thereafter, for any test point $\tilde{x}$, our sign prediction is simply ${\mathrm{sign}}((\tilde{x}W_{1}^{\star})_{+}W_{2}^{\star})$, which will have the same value as ${\mathrm{sign}}(\sum_{i}(\tilde{x}^{T}u_{i}^{\star})_{+}-(\tilde{x}^{T}u_{i}^{\prime\star})_{+}).$
We emphasize that our reframing of training a ReLU network for binary classification as linear programming does not eliminate the nonlinearity of the ReLU activation. Instead, this approach works because the ReLU activation patterns for a given training dataset are finite. By looping through these activation patterns, we can explicitly enumerate them. At test time, the ReLU nonlinearity is preserved, as our prediction ${\mathrm{sign}}(\sum_{i}(\tilde{x}^{T}u_{i})_{+}-(\tilde{x}^{T}u_{i}^{\prime})_{+})$ depends on the sign of $\tilde{x}^{T}u_{i}$ and $\tilde{x}^{T}u_{i}^{\prime}$, ensuring that the expressiveness of the nonlinearity remains intact. A careful reader might note that the number of patterns $P$ increases with the size of the training data, meaning the number of variables in Problem (3) may also grow. Additionally, finding all activation patterns poses a challenge. In Section 8, we demonstrate that subsampling a set of non-duplicate activation patterns performs well in practice. For further grounding, Appendix D.4 outlines an iterative hyperplane filtering method that guarantees the identification of all activation patterns with a reasonable complexity bound.
@@ 6 Training Deep Neural Networks via Cutting-Planes
We present our cutting-plane-based NN training algorithm here in main text. For generalizations, such as (i) relaxing the data distribution to remove the linear separability requirement (Appendix C.1), and (ii) extending from classification to regression (Appendix C.2), we refer readers to Appendix C. We emphasize that both the relaxed data constraint and the ability to handle regression tasks are unique to our method and have not been achieved in prior work.
@@
After we get the final $\theta_{c}^{t}$, for any test point $\tilde{x}$, we can directly compute our sign prediction with $\theta_{c}^{t}$. For example, for two-layer ReLU model, $\theta_{c}^{t}$ will be of form $\{u_{i}^{t},u_{i}^{\prime t}\}$, our final sign prediction would simply be ${\mathrm{sign}}(\sum_{i}(\tilde{x}^{T}u_{i}^{t})_{+}-(\tilde{x}^{T}u_{i}^{\prime t})_{+}).$ For deeper models, the prediction is a bit more complex and is given in equation (17) in Appendix D.2. Notably, the computation of final sign prediction from value of $\theta_{c}^{t}$ always takes a single step, just as one forward pass of original NN model formulation. Moreover, one can also restore optimal NN weights from final $\theta_{c}^{t}$, see Appendix D.1 for two-layer case reconstruction of optimal NN parameters and Appendix D.2 for general deep NN models.
@@
  • Center. The “center” function calculates the center of the convex set $\mathcal{T}^{t}$. There are a couple of notions of centers, such as center of gravity (CG), center of maximum volume ellipsoid (MVE), Chebyshev’s center, and analytic center (Boyd & Vandenberghe (2004)). Among these, the analytic center is the easiest to compute and is empirically known to be effective (Goffin et al. (1997), Atlason et al. (2008)). This is the notion of center that we will adopt to compute the query point in our algorithm. See Appendix F.1 for details.
@@ 7.1 Algorithm II: Cutting-Plane Localization for Active Learning
Our proposed cutting-plane-based active learning algorithm adapts and extends the generic framework discussed in Section 6 (Algorithm 3). For the sake of simplicity, we present in this section the algorithm specifically for binary classification and with respect to two-layer ReLU NNs. We emphasize that the algorithm can be easily adapted to the case of multi-class and regression, per discussions in Appendix C.2, and for deeper NNs following our reformulation in Theorem 5.2.
@@
For brevity, we present the algorithm for the first setup here and refer the rest to Appendix B. Algorithm 2 summarizes the proposed algorithm under the first setup, where we have used $\mathcal{D}_{\text{QS}}$ to denote the query synthesis. We note that Algorithm 4 for the second setup is obtained simply by changing $\mathcal{D}_{\text{QS}}$ to $\mathcal{D}_{\text{LQ}}$, which denotes limited query.
The general workflow of Algorithm 2 follows that of Algorithm 3 but on the transformed parameter space via the ReLU networks. We highlight here two key differences in our algorithm: (1) in each iteration, for faster empirical convergence, we query twice for the data point that is classified positively and the one that is classified negatively with highest confidence, respectively. If one or both of them turn out to be miss-classifications, this informs the active learner well and we expect a large cut. (2) For two- and three-layer ReLU networks, Ergen & Pilanci (2021b) demonstrate that these models can be reformulated as exact convex programs. This allows the option of incorporating a final convex solver into Algorithm 2, applied after the active learning loop with the data collected thus far. This convex reformulation includes regularization, which can improve the performance of our cutting-plane AL in certain tasks. For more details, see Appendix F.2
@@ 7.2 Convergence Guarantees
We give theoretical examination of the convergence properties of Algorithm 2 and 4 with respect to both the center of gravity (cg) and the center of maximum volume ellipsoid (MVE). Analysis of Algorithm 5 for inexact cuts is given in Appendix E.2. As the analysis of MVE closely parallels that of CG, we refer readers to Appendix E.3 for a detailed discussion, in the interest of brevity. For both centers, we measure the convergence speed with respect to the volume of the localization set $\mathcal{T}^{t}$ and judge the progress in iteration $t$ by the fractional decrease in volume: $\textbf{vol}(\mathcal{T}^{t+1})/\textbf{vol}(\mathcal{T}^{t}).$
@@ 8 Experiments
We validate our proposed training and active learning methods through extensive experiments, comparing them with various popular baselines from scikit-activeml (Kottke et al., 2021) and DeepAL (Huang, 2021). Synthetic data experiments are presented in Section 8.1, and real data experiments are presented in Section 8.2. An overview of each baseline is given in Appendix F.3. For implementation details and additional results, refer to Appendix F.
Figure 2: Decision boundaries for binary classification on the spiral dataset for the cutting-plane AL method using a two-layer ReLU neural network, alongside various deep AL baselines. For compactness, we also include the decision boundaries for the cutting-plane AL method with a three-layer ReLU network in the collage to demonstrate its feasibility. For fairness of comparison, we use the same two-layer ReLU network structure and embedding size of 623 for all methods. We enforce the same hyperparameters for all deep AL baselines and select the best performing number of training epochs at $2000$ and a learning rate at $0.001$ to ensure optimal performance. See Appendix F.3 for details.
@@ 8.1 Synthetic Data Experiments
We use a synthetic dataset of two intertwined spirals with positive and negative labels, respectively (see generation details in Appendix F.4). We generate 100 spiral data points with a 4:1 train-test split. Table 3 in Appendix F.4 presents the train and test accuracy of our cutting-plane AL (Algorithm 4) compared to popular deep AL baselines, with all methods evaluated with a query budget of 20 points (25% of train data). Our method achieved perfect accuracy on both sets, outperforming all baselines. Notably, the strong performance of our cutting-plane AL extends to the 3-layer case, achieving train/test accuracies of 0.71/0.60, while using only a fraction of neurons per layer (57 and 34, resp.) compared to the two-layer cutting-plane AL, which used 623 neurons. This result is illustrated in the corresponding decision boundary plot in Figure 2, where the 3-layer cutting-plane AL is one of the few methods to capture the spiral’s rough shape despite using smaller embeddings, while the two-layer cutting-plane AL, with the same network structure as the baselines, precisely traces the spiral. We emphasize that the superior performance of our cutting-plane AL remains consistent across different random seeds. As shown in the error-bar plot in Figure 15, our approach reliably converges to the optimal classifier faster than all the tested baselines in the number of queries.
@@
We evaluate the performance of our method (Algorithm 6) against the same seven baselines used in the spiral task, along with two additional popular regression AL methods from scikit-activeml: greedy sampling in target space (GreedyT) and KL divergence maximization (kldiv), on a simple quadratic regression task. We generate 100 noise-free data points from the function $y=x^{2}$ and apply the AL methods on a 4:1 train-test split with a query budget of 20 points. Table 5 in Appendix F.5 shows the root mean square error (RMSE) (see Definition 5), and left of Figure 3 provides a visualization of the final predictions of our method compared with a selection of baselines. We refer readers to Figure 18 for the complete version. Our cutting-plane AL achieves the lowest train/test RMSE of 0.01/0.01, representing an over 80%/75% reduction compared to the next best-performing baseline.This superior performance remains consistent across random seeds, as demonstrated in the train/test error-bar plots right of Figure 3, where our method consistently converges faster to the optimal classifier in terms of the number of queries.
Figure 3: Left: Predictions for the quadratic regression task using the cutting-plane AL method with a two-layer ReLU neural network, alongside various deep AL baselines. For brevity, we present only the most representative examples here and refer to Figure 18 in Appendix F.5 for the full result. Implementation details can be found in Appendix F.5. We note that the linear cutting-plane AL for regression (Algorithm 7) becomes infeasible when solving for the next center after the fourth query. This failure is expected, see Appendix B. Thus the reported prediction for this method is based on 4 queries, while all other methods use 20 queries. Right: Logarithm of mean test/train RMSE across seeds (0-4) versus the number of queries for the two-layer cutting-plane AL and various baselines. The linear cutting-plane method is omitted from the plot due to its infeasibilty.
@@ 8.2 Sentiment Classification for Real Datasets Using LLM Embeddings
To demonstrate the viability of our method when applied to tackle real life tasks, we also explore the concatenation of microsoft Phi-2 (Javaheripi et al., 2023) model with our two-layer ReLU binary classifier for sentiment classification task on IMDB (Maas et al., 2011) movie review datasets. Specifically, the dataset contains $50k$ movie reviews collected online with each comment accompanied with binary labels where $+1$ denoting a positive review and $-1$ denoting a negative review. See Appendix F.6 for several training data examples. We test our proposed active learning algorithm conjuncted with our cutting-plane training scheme as described in Section 6 and Section 7 respectively. For our experiment, we first collect all last layer Phi-2 embeddings (corresponding to last token), which is of size $d=2560$, for our training and testing movie reviews as our feature vectors. We follow the implementation details in Appendix F.1 and sample $P=500$ activation patterns to cater for the large dimension feature space.
@@ 9 Conclusion and Limitation
Despite its novelty, our current implementation has several key limitations that hinder its competitiveness with large-scale models trained using gradient-based methods. First, although we employ subsampling of activation patterns and propose an iterative filtering scheme (see Appendix D.4), the subsampling process is not exhaustive, which impacts model performance, especially with high-dimensional data. Refining the activation pattern sampling strategy could significantly improve results. Second, we rely on analytic center retrieval during training, which we solve using CVXPY. However, this solution is CPU-bound and becomes inefficient for large-scale problems with many variables. Developing a center-finding algorithm that leverages GPU parallelism is crucial to unlocking the full potential of our training method. Finally, while current large language model (LLM) training often involves cross-entropy loss, our approach has so far been applied only to classification and regression tasks. Extending our method to handle more diverse loss functions presents an exciting avenue for future research.

### v3

@@ 4 Key Observation: Training ReLU NNs for Binary Classification is Linear Programming
To break this bottleneck and extend the cutting-plane-based learning method to nonlinear models, we observe that training a ReLU network for binary classification can be formulated as a linear programming problem. This insight is crucial for extending the learning scheme in (Louche & Ralaivola, 2015) to more complex models. We now develop our core idea of reframing binary classification with ReLU models as linear programs. For clarity, we present our results in two theorems: one for two-layer ReLU networks and another for ReLU networks of arbitrary depth. We focus on the two-layer case in the main paper for detailed discussion, deferring the more abstract general case to Appendix E.2. Since the general case is an extension of the two-layer model, focusing on the two-layer case should provide a clearer understanding of the core concepts.
@@
Therefore, if we iterate over all $D_{i}\in\mathcal{D}$, we are guaranteed to reach each ${K}_{i}$. The equivalence in Theorem 4.1 holds in the sense that, which we prove rigorously in Appendix E.1, whenever there is solution $W_{1},W_{2}$ to Problem (2), there is always solution $\{u_{i},u_{i}^{\prime}\}$ to problem (3) and vice versa. Moreover, when an optimal solution $\{u_{i}^{\star},u_{i}^{\prime\star}\}$ to problem (3) has been found, we can explicitly create an optimal solution $\{W_{1}^{\star},W_{2}^{\star}\}$ to Problem (2) from value of $\{u_{i}^{\star},u_{i}^{\prime\star}\}$, see Appendix E.1 for details. Thereafter, for any test point $\tilde{x}$, our sign prediction is simply ${\mathrm{sign}}((\tilde{x}W_{1}^{\star})_{+}W_{2}^{\star})$, which will have the same value as ${\mathrm{sign}}(\sum_{i}(\tilde{x}^{T}u_{i}^{\star})_{+}-(\tilde{x}^{T}u_{i}^{\prime\star})_{+}).$
We emphasize that our reframing of training a ReLU network for binary classification as linear programming does not eliminate the nonlinearity of the ReLU activation. Instead, this approach works because the ReLU activation patterns for a given training dataset are finite. By looping through these activation patterns, we can explicitly enumerate them. At test time, the ReLU nonlinearity is preserved, as our prediction ${\mathrm{sign}}(\sum_{i}(\tilde{x}^{T}u_{i})_{+}-(\tilde{x}^{T}u_{i}^{\prime})_{+})$ depends on the sign of $\tilde{x}^{T}u_{i}$ and $\tilde{x}^{T}u_{i}^{\prime}$, ensuring that the expressiveness of the nonlinearity remains intact. A careful reader might note that the number of patterns $P$ increases with the size of the training data, meaning the number of variables in Problem (3) may also grow. Additionally, finding all activation patterns poses a challenge. In Section 7, we demonstrate that subsampling a set of non-duplicate activation patterns performs well in practice. For further grounding, Appendix E.4 outlines an iterative hyperplane filtering method that guarantees the identification of all activation patterns with a reasonable complexity bound.
@@ 5 Training Deep Neural Networks via Cutting-Planes
We present our cutting-plane-based NN training algorithm here in main text. For generalizations, such as (i) relaxing the data distribution to remove the linear separability requirement (Appendix D.1), and (ii) extending from classification to regression (Appendix D.2), we refer readers to Appendix D. We emphasize that both the relaxed data constraint and the ability to handle regression tasks are unique to our method and have not been achieved in prior work.
@@
After we get the final $\theta_{c}^{t}$, for any test point $\tilde{x}$, we can directly compute our sign prediction with $\theta_{c}^{t}$. For example, for two-layer ReLU model, $\theta_{c}^{t}$ will be of form $\{u_{i}^{t},u_{i}^{\prime t}\}$, our final sign prediction would simply be ${\mathrm{sign}}(\sum_{i}(\tilde{x}^{T}u_{i}^{t})_{+}-(\tilde{x}^{T}u_{i}^{\prime t})_{+}).$ For deeper models, the prediction is a bit more complex and is given in equation (17) in Appendix E.2. Notably, the computation of final sign prediction from value of $\theta_{c}^{t}$ always takes a single step, just as one forward pass of original NN model formulation. Moreover, one can also restore optimal NN weights from final $\theta_{c}^{t}$, see Appendix E.1 for two-layer case reconstruction of optimal NN parameters and Appendix E.2 for general deep NN models.
@@
  • Center. The “center” function calculates the center of the convex set $\mathcal{T}^{t}$. There are a couple of notions of centers, such as center of gravity (CG), center of maximum volume ellipsoid (MVE), Chebyshev’s center, and analytic center (Boyd & Vandenberghe (2004)). Among these, the analytic center is the easiest to compute and is empirically known to be effective (Goffin et al. (1997), Atlason et al. (2008)). This is the notion of center that we will adopt to compute the query point in our algorithm. See Appendix G.1 for details.
@@ 6.1 Algorithm II: Cutting-Plane Localization for Active Learning
Our proposed cutting-plane-based active learning algorithm adapts and extends the generic framework discussed in Section 5 (Algorithm 3). For the sake of simplicity, we present in this section the algorithm specifically for binary classification and with respect to two-layer ReLU NNs. We emphasize that the algorithm can be easily adapted to the case of multi-class and regression, per discussions in Appendix D.2, and for deeper NNs following our reformulation in Theorem 4.2.
@@
For brevity, we present the algorithm for the first setup here and refer the rest to Appendix C. Algorithm 2 summarizes the proposed algorithm under the first setup, where we have used $\mathcal{D}_{\text{QS}}$ to denote the query synthesis. We note that Algorithm 4 for the second setup is obtained simply by changing $\mathcal{D}_{\text{QS}}$ to $\mathcal{D}_{\text{LQ}}$, which denotes limited query.
The general workflow of Algorithm 2 follows that of Algorithm 3 but on the transformed parameter space via the ReLU networks. We highlight here two key differences in our algorithm: (1) in each iteration, for faster empirical convergence, we query twice for the data point that is classified positively and the one that is classified negatively with highest confidence, respectively. If one or both of them turn out to be miss-classifications, this informs the active learner well and we expect a large cut. (2) For two- and three-layer ReLU networks, Ergen & Pilanci (2021b) demonstrate that these models can be reformulated as exact convex programs. This allows the option of incorporating a final convex solver into Algorithm 2, applied after the active learning loop with the data collected thus far. This convex reformulation includes regularization, which can improve the performance of our cutting-plane AL in certain tasks. For more details, see Appendix G.2
@@ 6.2 Convergence Guarantees
We give theoretical examination of the convergence properties of Algorithm 2 and 4 with respect to both the center of gravity (cg) and the center of maximum volume ellipsoid (MVE). Analysis of Algorithm 5 for inexact cuts is given in Appendix F.2. As the analysis of MVE closely parallels that of CG, we refer readers to Appendix F.3 for a detailed discussion, in the interest of brevity. For both centers, we measure the convergence speed with respect to the volume of the localization set $\mathcal{T}^{t}$ and judge the progress in iteration $t$ by the fractional decrease in volume: $\textbf{vol}(\mathcal{T}^{t+1})/\textbf{vol}(\mathcal{T}^{t}).$
@@ 7 Experiments
We validate our proposed training and active learning methods through extensive experiments, comparing them with various popular baselines from scikit-activeml (Kottke et al., 2021) and DeepAL (Huang, 2021). Synthetic data experiments are presented in Section 7.1, and real data experiments are presented in Section 7.2. An overview of each baseline is given in Appendix G.3. For implementation details and additional results, refer to Appendix G.
Figure 2: Decision boundaries for binary classification on the spiral dataset for the cutting-plane AL method using a two-layer ReLU neural network, alongside various deep AL baselines. For compactness, we also include the decision boundaries for the cutting-plane AL method with a three-layer ReLU network in the collage to demonstrate its feasibility. For fairness of comparison, we use the same two-layer ReLU network structure and embedding size of 623 for all methods. We enforce the same hyperparameters for all deep AL baselines and select the best performing number of training epochs at $2000$ and a learning rate at $0.001$ to ensure optimal performance. See Appendix G.3 for details.
@@ 7.1 Synthetic Data Experiments
We use a synthetic dataset of two intertwined spirals with positive and negative labels, respectively (see generation details in Appendix G.4). We generate 100 spiral data points with a 4:1 train-test split. Table 3 in Appendix G.4 presents the train and test accuracy of our cutting-plane AL (Algorithm 4) compared to popular deep AL baselines, with all methods evaluated with a query budget of 20 points (25% of train data). Our method achieved perfect accuracy on both sets, outperforming all baselines. Notably, the strong performance of our cutting-plane AL extends to the 3-layer case, achieving train/test accuracies of 0.71/0.60, while using only a fraction of neurons per layer (57 and 34, resp.) compared to the two-layer cutting-plane AL, which used 623 neurons. This result is illustrated in the corresponding decision boundary plot in Figure 2, where the 3-layer cutting-plane AL is one of the few methods to capture the spiral’s rough shape despite using smaller embeddings, while the two-layer cutting-plane AL, with the same network structure as the baselines, precisely traces the spiral. We emphasize that the superior performance of our cutting-plane AL remains consistent across different random seeds. As shown in the error-bar plot in Figure 15, our approach reliably converges to the optimal classifier faster than all the tested baselines in the number of queries.
@@
We evaluate the performance of our method (Algorithm 6) against the same seven baselines used in the spiral task, along with two additional popular regression AL methods from scikit-activeml: greedy sampling in target space (GreedyT) and KL divergence maximization (kldiv), on a simple quadratic regression task. We generate 100 noise-free data points from the function $y=x^{2}$ and apply the AL methods on a 4:1 train-test split with a query budget of 20 points. Table 5 in Appendix G.5 shows the root mean square error (RMSE) (see Definition 5), and left of Figure 3 provides a visualization of the final predictions of our method compared with a selection of baselines. We refer readers to Figure 18 for the complete version. Our cutting-plane AL achieves the lowest train/test RMSE of 0.01/0.01, representing an over 80%/75% reduction compared to the next best-performing baseline.This superior performance remains consistent across random seeds, as demonstrated in the train/test error-bar plots right of Figure 3, where our method consistently converges faster to the optimal classifier in terms of the number of queries.
Figure 3: Left: Predictions for the quadratic regression task using the cutting-plane AL method with a two-layer ReLU neural network, alongside various deep AL baselines. For brevity, we present only the most representative examples here and refer to Figure 18 in Appendix G.5 for the full result. Implementation details can be found in Appendix G.5. We note that the linear cutting-plane AL for regression (Algorithm 7) becomes infeasible when solving for the next center after the fourth query. This failure is expected, see Appendix C. Thus the reported prediction for this method is based on 4 queries, while all other methods use 20 queries. Right: Logarithm of mean test/train RMSE across seeds (0-4) versus the number of queries for the two-layer cutting-plane AL and various baselines. The linear cutting-plane method is omitted from the plot due to its infeasibilty.
@@ 7.2 Sentiment Classification for Real Datasets Using LLM Embeddings
To demonstrate the viability of our method when applied to tackle real life tasks, we also explore the concatenation of microsoft Phi-2 (Javaheripi et al., 2023) model with our two-layer ReLU binary classifier for sentiment classification task on IMDB (Maas et al., 2011) movie review datasets. Specifically, the dataset contains $50k$ movie reviews collected online with each comment accompanied with binary labels where $+1$ denoting a positive review and $-1$ denoting a negative review. See Appendix G.6 for several training data examples. We test our proposed active learning algorithm conjuncted with our cutting-plane training scheme as described in Section 5 and Section 6 respectively. For our experiment, we first collect all last layer Phi-2 embeddings (corresponding to last token), which is of size $d=2560$, for our training and testing movie reviews as our feature vectors. We follow the implementation details in Appendix G.1 and sample $P=500$ activation patterns to cater for the large dimension feature space.
@@ 8 Conclusion and Limitations
Despite its novelty, our current implementation has several key limitations that hinder its competitiveness with large-scale models trained using gradient-based methods. First, although we employ subsampling of activation patterns and propose an iterative filtering scheme (see Appendix E.4), the subsampling process is not exhaustive, which impacts model performance, especially with high-dimensional data. Refining the activation pattern sampling strategy could significantly improve results. Second, we rely on analytic center retrieval during training, which we solve using CVXPY. However, this solution is CPU-bound and becomes inefficient for large-scale problems with many variables. Developing a center-finding algorithm that leverages GPU parallelism is crucial to unlocking the full potential of our training method. Finally, while current large language model (LLM) training often involves cross-entropy loss, our approach has so far been applied only to classification and regression tasks. Extending our method to handle more diverse loss functions presents an exciting avenue for future research.